Question 1 1 pts You invested a sum of money 6 years ago at 10% simple interest. Today the investment is worth $18,000. How much was your original investment? Express your answer in $to the nearest whole $.
Answer #1

Following is the formula of future value in the case of simple interest.

Future value = P(1 + rt)

In which, P is the principal amount, r is the rate of interest (here, r = 10% = 0.10) and t is the number of years (here, t =6)

So, putting the values, we get:

18,000 = P(1 + 0.10*6)

or, 18,000 = 1.6P

or, P = 18,000/1.6 = $11,250.

So, the original investment was $11,250.